OpenStudy (anonymous):

Evaluate the limits. Use -∞ or ∞ where appropriate. Evaluate lim (x^2-3x+8) as x approaches 2

OpenStudy (anonymous):

lim┬(x→2)⁡〖(x^2- 3x+8)^ 〗

OpenStudy (abb0t):

Just plug in 2.

OpenStudy (anonymous):

the limit as x approaches 2 is 6.

OpenStudy (abb0t):

How did you get 6?

OpenStudy (anonymous):

2^2 = 4 4-6 = -2 -2+8=6

OpenStudy (abb0t):

good job.

OpenStudy (anonymous):

So is that correct? or was I suppose to factor the equation first?

OpenStudy (abb0t):

No, that is correct.

OpenStudy (anonymous):

Thanks for the help